OpenStudy (anonymous):

In the expansion of (x+y)^15, a certain term contains x^10. What is the exponent of y in this term? a. 8 b. 9 c. 11 d. 5

OpenStudy (asnaseer):

when you expand (x+y)^15 the first term will have just x^15 and the last term will have just y^15. In between these you will get terms involving both x and y. The thing that remains the same for each term is the total of the powers of the x and y in each term will always be 15.

myininaya (myininaya):

So each term in the expansion will look like this \[c x^ay^b \text{ where } a+b=15 \text{ and c is some constant} \]

OpenStudy (anonymous):

@Lulu49 are you okie from here?

OpenStudy (anonymous):

Still a bit confused. is the answer 11 or 5?

OpenStudy (anonymous):

Total of the powers is 15 You already known the power of x is 10 -> Power of y = ...

myininaya (myininaya):

If a=10 , then how do you solve 10+b=15 for b?

OpenStudy (anonymous):

I think the answer is 5 because it goes into both 10 and 15

OpenStudy (anonymous):

:)

myininaya (myininaya):

the answer is 5 because 10+5=15

OpenStudy (anonymous):

Oh alright thank you both (: